OpenStudy (anonymous):

How do you know that a Quadratic equation cannot be solved by factoring, but can be solved by using the quadratic formula?

jimthompson5910 (jim_thompson5910):

If the discriminant isn't a perfect square, then you cannot solve the equation by factoring.

myininaya (myininaya):

well really any quadratic equation can be solved by factoring or by quadratic formula

jimthompson5910 (jim_thompson5910):

oh I meant factoring over the rationals

myininaya (myininaya):

i have proved this

myininaya (myininaya):

although it is hard to factor over the complex it is not impossible

OpenStudy (anonymous):

Usually I try by factoring method if it fails then I resort to quadratic formula which works for ANY quadratic equation.

jimthompson5910 (jim_thompson5910):

by "factor" I think they mean do so without the quadratic formula (ie find two numbers that multiply to this and add to that...etc)

myininaya (myininaya):

usually people just try to factor over the integers then if that is not possible they result to the quadratic formula (or completing the square)

myininaya (myininaya):

i can factor without the quadratic formula

jimthompson5910 (jim_thompson5910):

lol well average people then

myininaya (myininaya):

give me any quadratic and i will show you

OpenStudy (anonymous):

For instance: x^2 + 3x + 5 this is not factorable so you have to use the quadratic formula and you will get a complex root.

jimthompson5910 (jim_thompson5910):

not superhuman machines lol

OpenStudy (anonymous):

lol

OpenStudy (zarkon):

\[\left(x-\frac{-3}{2}-\frac{\sqrt{11}}{2}i\right)\left(x-\frac{-3}{2}+\frac{\sqrt{11}}{2}i\right)\]

myininaya (myininaya):

we can use factor by grouping as i did in this file above

OpenStudy (anonymous):

0.0 wow

OpenStudy (anonymous):

cool!!...learnt something new today

myininaya (myininaya):

i come up with that way

myininaya (myininaya):

:)

myininaya (myininaya):

i made a proof for the quadratic formula

myininaya (myininaya):

but its run time is much longer than completing the square

myininaya (myininaya):

so it almost useless

OpenStudy (anonymous):

thank you for all the help ^.^